# Solving the Poisson equation using coupled Markov chains

### v3

@@ 5.2 Hamiltonian Monte Carlo for Bayesian logistic regression
5.3 Gibbs sampler for linear regression with shrinkage prior
We examine a more challenging example, with a Bayesian linear regression of $n=71$ responses on $p=4088$ predictors of the riboflavin data set (Bühlmann et al., 2014), using a continuous shrinkage prior on the coefficients (e.g. Bhadra et al., 2019). The model, MCMC algorithm and its coupling are taken from Biswas et al. (2022) and a self-contained description is provided in Appendix G. Computational difficulties arise from the use of the shrinkage prior, that induces multimodality and heavy tails in the posterior distribution of the coefficients, denoted by $\beta$. The target is defined on the space of coefficients, global precision, local precisions, and the variance of the observation noise. The state space is of dimension $81782p+2=$ 8178. In this example, the meeting times have been shown to have Exponential tails in Biswas et al. (2022, Proposition 6) under the assumption that the global precision $\xi$ has a compact support, in which case Assumption 2 holds for any $\kappa$. In the experiments, we do not restrict the support of $\xi$, and we use a half-Cauchy distribution on $\xi^{-1/2}$.
Figure 5.5: Gibbs sampler for linear regression with shrinkage prior. Left: trace of the component $\beta_{2564}$ of three independent chains. Middle: histogram of $\beta_{2564}$, obtained from long MCMC runs. Right: upper bounds on $|\pi_{0}P^{t}-\pi|_{\text{TV}}$.
Based on preliminary runs, we choose the test function $h:x\mapsto\beta_{2564}$, which is a coordinate of the regression coefficients with a clearly bimodal marginal posterior distribution. Figure 5.5(a) shows three independent traces of $\beta_{2564}$ over 1000 iterations of the chain. Figure 5.5(b) presents a histogram of $\beta_{2564}$, obtained from 10 independent chains run for 50,000 iterations each and discarding 2000 iterations as burn-in. Figure 5.5(c) shows upper bounds on $|\pi_{0}P^{t}-\pi|_{\text{TV}}$ obtained with the method of Biswas et al. (2019), using a lag $L=1000$ and $10^{3}$ independent meeting times. From this we choose $k=L=1000$ and $\ell=5k$.
To define $g=g_{y}$, we draw $y$ once from the prior, and keep it fixed. We generate $M=10^{3}$ independent estimates of $v(P,h)$, for $R\in\{1,5,10\}$. The results are summarized in Table 5.4. We again observe tangible gains in efficiency when increasing $R$, with diminishing returns. Overall we obtain relatively precise information about $v(P,h)$.
R estimate total cost fishy cost variance of estimator inefficiency
1 [77 - 98] [12308 - 12383] [1522 - 1594] [2.2e+04 - 3.2e+04] [2.7e+08 - 4.1e+08]
5 [78 - 87] [18469 - 18635] [7683 - 7836] [5.4e+03 - 6.9e+03] [1e+08 - 1.3e+08]
10 [78 - 85] [26219 - 26444] [15428 - 15652] [2.6e+03 - 3.1e+03] [6.7e+07 - 8.1e+07]
Table 5.4: Gibbs sampler for linear regression with shrinkage prior: proposed estimators of $v(P,h)$.
Finally, we compare with batch means and spectral variance estimators, as in the previous section. We compute BM and SV from $M=25$ independent runs, each of which involves $4$ chains. We discard the first 1,500 discarded as burn-in, and run the chains for $t=2\cdot 10^{5}$ iterations. Figure 5.6 shows the resulting estimates. The BM and SV estimates have a low variability compared to UPAVE. For example, the estimators “SV r = 2” with $t=2\cdot 10^{5}$ have a cost of $8\times 10^{5}$ and a variance of $4$. To match that variance, one would need to average about 700 UPAVE runs with $R=10$, which would cost about $2\times 10^{7}$ units of transitions, that is 23 times more than “SV r = 2”. Meanwhile, the bias of batch means and spectral variance is noticeable in Figure 5.6.
Figure 5.6: Gibbs sampler for linear regression with shrinkage prior: batch mean and spectral variance estimators, from runs of length $2\cdot 10^{5}$. Each dot represents an estimate of $v(P,h)$. The yellow triangles represent the means for each method. The horizontal lines represent the estimate of $v(P,h)$ and a $95\%$ confidence interval from 1000 runs of UPAVE with $R=10$.
5.4 Particle marginal Metropolis–Hastings
We consider the state space model (SSM) example in Middleton et al. (Section 4.2 2020), inspired by a model capturing the activation of neuron of rats when responding to a periodic stimulus. The observations are counts of neuron activations over $50$ experiments. We consider $100$ data points represented in Figure 5.7(a). They are modelled as
$y_{t}|x_{t}\sim\text{Binomial}(50,\text{logistic}(x_{t})),$
where $\text{logistic}:x\mapsto 1/(1+\exp(-x))$ and
The prior is $\text{Uniform}(0,1)$ on $\alpha$, and $\sigma^{2}$ is fixed to 1.5 here for simplicity. The likelihood is intractable but can be estimated using a particle filter. As in Middleton et al. (2020) we use controlled SMC (Heng et al., 2020), and we plug the likelihood estimator in the particle marginal Metropolis–Hastings algorithms (PMMH, Andrieu et al., 2010). We use 3 iterations of controlled SMC at each PMMH iteration. The proposal on $\alpha$ is a Normal random walk, with a standard deviation drawn from $\text{Uniform}(0.001,0.2)$ at each iteration. The coupling operates with a reflection-maximal coupling of the proposals, and independent runs of SMC if the proposals differ. Verification of Assumption 2 for the algorithm employed here is an open question. Relevant comments can be found in Middleton et al. (2020, Section 2.3). We initialize the chains from the prior $\text{Uniform}(0,1)$. An approximation of the posterior distribution is shown in Figure 5.7(b), obtained from 100 chains of length $20,000$ and a burn-in of $1000$ steps.
Figure 5.7: Particle marginal Metropolis–Hastings: 100 observations (left) and posterior distribution on $\alpha$ (right) approximated with particle marginal Metropolis–Hastings.
Expecting PMMH to be polynomially ergodic, we examine the tails of the distribution of the meeting times. We generate $10^{5}$ meeting times, either using 64 or 256 particles in each run of SMC within PMMH. The empirical survival functions of the meeting times $\tau$, or more exactly of $\tau-L$ with a lag $L=100$, are shown in Figure 5.8. Since both axes are on logarithmic scale, a straight line indicates a polynomial decay for $\mathbb{P}(\tau>t)$. We indeed observe straight lines on the parts of figure where $t$ is large enough. Using linear regression we estimate the polynomial rate to be around 1 when using 64 particles (focusing on $t>200$), and above 2 when using 256 particles (focusing on $t>100$).
Figure 5.8: Particle marginal Metropolis–Hastings: survival function $\mathbb{P}(\tau>t)$ when using 64 particles (left) or 256 particles (right) in PMMH. Both axes use logarithmic scale.
Given the heavy tails of $\tau$ when using 64 particles, we were not able to reliably estimate the associated $v(P,h)$. We thus focus on the use of 256 particles, and we generate UPAVE to estimate $v(P,h)$, $M=500$ times independently, for $h:x\mapsto x$. We set $y=0.5$ in the definition of the fishy function estimator $G_{y}$. We use $k=L=500$ and $\ell=5k$ for unbiased MCMC approximations. We choose $R=50$, the number of atoms at which $g_{y}$ is estimated per signed measure. From the UPAVE runs, we can extract all the locations at which $g_{y}$ is estimated by $G_{y}$, along with the estimates. We then represent an approximation of $g_{y}$ in Figure 5.9(a), and a histogram of the $500$ estimates of $v(P,h)$ in Figure 5.9(b). We see that the relative variance is fairly large, and notice that many estimates are negative.
Figure 5.9: Particle marginal Metropolis–Hastings: using $y=0.5$, estimation of $g_{y}(x)$ (left) and histogram of proposed estimators of $v(P,h)$ (right).
We then change $y$ from $0.5$ to $0.975$, i.e. we place $y$ in the middle of the posterior distribution as shown in Figure 5.7(b), and reproduce the same plots in Figure 5.10. We see that the fishy function takes smaller values and its estimation is more precise. As a result, the distribution of $\hat{v}(P,h)$ is considerably more concentrated. The effect of the choice of $y$ is summarized in Table 5.5, where all entries are confidence intervals based on the nonparametric bootstrap. We observe that the choice of $y$ impacts the cost of fishy function estimation, as well as its variance and thus the variance of UPAVE. Here this results in orders of magnitude of difference in efficiency.
Figure 5.10: Particle marginal Metropolis–Hastings: using $y=0.975$, estimation of $g_{y}(x)$ (left) and histogram of proposed estimators of $v(P,h)$ (right).
y estimate total cost fishy cost variance of estimator
Table 5.5: Particle marginal Metropolis–Hastings: effect of $y$ on the proposed confidence interval for $v(P,h)$, the cost of fishy function estimation, the variance of $\hat{v}(P,h)$ and its inefficiency, based on $M=500$ independent repeats, and using $R=50$, $k=L=500$, $\ell=5k$.
Figure 5.11 presents 25 BM and SV estimates, each based on 4 parallel chains. The chains are of length $40,000$ and the first $1000$ iterations are discarded as burn-in. Here the bias is hard to notice. Meanwhile, the variability of BM and SV is comparable with that of UPAVE: for example the SV estimator with $r=2$ has a variance of $3.3\cdot 10^{-8}$, comparable to that of an average of 18 UPAVE with $R=50$, which would amount to a cost of 110,000 transitions, which is less than the cost of producing the SV estimator with 4 chains of length $40,000$.
Figure 5.11: Particle marginal Metropolis–Hastings: batch mean and spectral variance estimators, from runs of length $4\cdot 10^{4}$. Each dot represents an estimate of $v(P,h)$. The yellow triangles represent the means for each method. The horizontal lines represent the estimate of $v(P,h)$ and a $95\%$ confidence interval obtained from 500 runs of UPAVE with $R=50$.
@@ 6 Discussion
As with other works on unbiased MCMC (Atchadé & Jacob, 2024), it is worth emphasizing that the performance depends on the underlying MCMC algorithm, its initialization and its coupling. We refer to the bimodal target in Section 5.1 of Jacob et al. (2020) for a situation where multimodality in the target distribution combined with a poor design of the MCMC sampler gives misleading estimates in finite samples, despite the lack of bias and a provably finite variance. In our theoretical results, we have prioritized assumptions on the moments of meeting times of the coupled Markov chains under strong but reasonable initialization assumptions, which can be cleanly separated from assumptions on the moments of functions. Given this emphasis, the results appear to be fairly strong and provide a sensible relationship between the moments of the meeting time and moments of the proposed estimators.
Appendix A provides reminders on couplings of MCMC algorithms. Appendix B contains the proofs of our results. Appendix C provides reminders on unbiased MCMC. Appendix D reviews batch means and spectral variance estimators of the long-run variance of Markov chain ergodic averages. Code to reproduce the figures of this article can be found at: https://github.com/pierrejacob/unbiasedpoisson/.
Acknowledgments.
Research of AL supported by EPSRC grant ‘CoSInES (COmputational Statistical INference for Engineering and Security)’ (EP/R034710/1). Research of DV supported by SERB (SPG/2021/001322).
Appendix A Coupling of MCMC algorithms
For completeness we recall some implementable coupling techniques for MCMC. We focus on Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithms (Hastings 1970). The goal is to couple generated trajectories such that exact meetings can occur. Relevant considerations can be found in Jacob et al. (2020), Wang et al. (2021), Atchadé & Jacob (2024).
@@ Appendix A Coupling of MCMC algorithms
When $\mu$ and $\nu$ are Normal distributions with the same variance, an alternative maximal coupling procedure is described in Algorithm A.3; it was proposed in Bou-Rabee et al. (2020). Its cost is deterministic and independent of $\mu$ and $\nu$. In the algorithmic description, $\varphi$ refers to the probability density function of the standard Normal distribution. We use Algorithm A.3 to couple the Metropolis-adjusted Langevin algorithm in Section 5.2 and Appendix F.
@@ B.1 Assumption on the meeting time
Applying Proposition 24 below to $\psi(v)=\phi(v)=\vartheta v^{\alpha}$, there exists a constant $\beta$ such that for all $x,y\in{\mathcal{X}}$ and all $n\in{\mathbb{N}}$,
@@
In particular, with this $\phi$ and any $\sigma\in(0,1)$ there exists $c_{\sigma}>0$ such that $H_{\phi}^{-1}(\sigma n)\geq c_{\sigma}H_{\phi}^{-1}(n)$ and concavity and non-negativity of $\phi$ allows us to conclude that $\phi\circ H_{\phi}^{-1}(\sigma n)\geq c_{\sigma}\,\phi\circ H_{\phi}^{-1}(n)$. Finally, there exists $c^{\prime}_{\vartheta,\alpha}>0$ such that $\phi\circ H_{\phi}^{-1}(n)\geq c^{\prime}_{\vartheta,\alpha}(1+n)^{\frac{\alpha}{1-\alpha}}$. According to Proposition 4.3.2 in Douc et al. (2018), the drift condition $PV(x)\leq V(x)-\vartheta V^{\alpha}(x)+b\mathds{1}(x\in\mathcal{C})$ implies $\pi(V^{\alpha})<\infty$, which allows to apply Proposition 8 and the proof is concluded. ∎
Proposition 24.
@@
$\int\pi({\rm d}x)\pi({\rm d}y){\mathbb{P}}_{x,y}(\tau>n)\leq\frac{\beta_{0}}{\phi\circ H_{\phi}^{-1}(n)},$
@@
Proof of Proposition 24.
In this proof, for any function $\phi$, we use the notation $H_{\phi}(t)=\int_{1}^{t}\frac{1}{\phi(u)}{\rm d}u$ and $r_{\phi}(t)=\phi\circ H_{\phi}^{-1}(t)$. By Lemma 19.5.3 in Douc et al. (2018), for any $\sigma\in(0,1-b/d)$, $\bar{P}\bar{V}\leq\bar{V}-\bar{\phi}\circ\bar{V}+\bar{b}\mathds{1}_{\bar{\mathcal{C}}}$ where $\bar{\phi}=\sigma\phi$, $\bar{b}=2b$, $\bar{V}(x,y)=V(x)+V(y)-1$.
@@
where $\bar{\mathcal{C}}={\mathcal{C}}\times{\mathcal{C}}$ and $b^{\prime}=\bar{b}r_{\phi}(1)/r_{\phi}^{2}(0)$. Set $r^{0}_{\bar{\phi}}(n)=\sum_{j=0}^{n}r_{\bar{\phi}}(j)$ and define $M=\sup_{k}-r_{\bar{\phi}}^{0}(k-1)+b^{\prime}r_{\bar{\phi}}(k)/\epsilon$, which is finite since $r_{\bar{\phi}}(k)/r^{0}_{\bar{\phi}}(k-1)\to 0$ as $k$ tends to infinity.
@@
where $\bar{\Delta}(x,y)=\mathds{1}(x\neq y)$. The coupling kernel $\bar{P}$ is sticky, i.e. $\bar{P}\bar{\Delta}(x,y)=0$ if $\bar{\Delta}(x,y)=0$. Hence, if $x=y$, the l.h.s. and r.h.s of (B.2) are null. We now assume that $x\neq y$. Using $\bar{\Delta}\leq 1$, we get
@@
where we have used that $b^{\prime}r_{\bar{\phi}}(k)-\epsilon(r^{0}_{\bar{\phi}}(k-1)+M)\leq 0$ by definition of $M$. Since $x\neq y$, we have $\bar{\Delta}(x,y)=1$ and hence $\bar{P}[\bar{\Delta}S_{k+1}](x,y)\leq[\bar{\Delta}S_{k}](x,y)$. Finally (B.2) is proved for all $x,y\in{\mathcal{C}}$. Now write
@@
$A_{n}=\left\{(x,y);\;V(x)+V(y)>H^{-1}_{\bar{\phi}}(n)\right\},$
@@
Proposition 25 (Douc et al. 2018, Proposition 21.2.3).
@@
and $P$ is aperiodic and ergodic of degree 2.
@@
and since Assumption 2 provides $\kappa>1$ and $\mathbb{E}_{x,\pi}[\tau^{\kappa}]<\infty$ for $\pi$-almost all $x$, $\lim_{t\to\infty}\left\|P^{t}(x,\cdot)-\pi\right\|_{{\rm TV}}=0$ for $\pi$-almost all $x$ and so $P$ is aperiodic by Douc et al. (2018, Lemma 9.3.9). We also observe that (B.4) implies (B.3) and since $\kappa>1$ we have
@@
$\leq\mathbb{E}_{x,\pi}\left[\mathbb{I}(\tau>t)\left\{\left|h(X_{t})\right|+\left|h(Y_{t})\right|\right\}\right],$
@@
$\left|P^{t}h(x)\right|^{p}$ $\leq\mathbb{E}_{x,\pi}\left[\mathbb{I}(\tau>t)\left\{\left|h(X_{t})\right|+\left|h(Y_{t})\right|\right\}\right]^{p}$
$\leq\mathbb{E}_{x,\pi}\left[\mathbb{I}(\tau>t)\left\{\left|h(X_{t})\right|+\left|h(Y_{t})\right|\right\}^{p}\right]$
$\leq 2^{p-1}\mathbb{E}_{x,\pi}\left[\mathbb{I}(\tau>t)\left\{\left|h(X_{t})\right|^{p}+\left|h(Y_{t})\right|^{p}\right\}\right].$
@@
$\leq 2^{p-1}\mathbb{E}_{\pi\otimes\pi}\left[\mathbb{I}(\tau>t)\left\{\left|h(X_{t})\right|^{p}+\left|h(Y_{t})\right|^{p}\right\}\right]$
@@
and we conclude by appealing to Proposition 25. ∎
Remark 28.

### v4

@@ 6 Discussion
As with other works on unbiased MCMC (Atchadé & Jacob, 2024), it is worth emphasizing that the performance depends on the underlying MCMC algorithm, its initialization and its coupling. We refer to the bimodal target in Section 5.1 of Jacob et al. (2020) for a situation where multimodality in the target distribution combined with a poor choice of MCMC sampler gives misleading estimates, despite the lack of bias and finite variance. In our theoretical results, we have prioritized assumptions on the moments of meeting times of the coupled Markov chains under strong but reasonable initialization assumptions, which can be cleanly separated from assumptions on the moments of functions. Given this emphasis, the results appear to be fairly strong and provide a sensible relationship between the moments of the meeting time and moments of the proposed estimators.
The appendices are structured as follows:
  • Appendix A
    provides reminders on unbiased MCMC with couplings.
  • Appendix B
    contains the proofs of our results.
  • Appendix C
    describes some existing results on the convergence of batch means and spectral variance estimators.
  • Appendix D
    verifies Assumption 2 quantitatively for the AR(1) process considered in Section 5.1.
  • Appendix E
    describes couplings of Hamiltonian Monte Carlo employed in Section 5.2.
  • Appendix F
    provides experiments with a Gibbs sampler for high-dimensional regression with shrinkage prior.
  • Appendix G
    provides experiments with a particle marginal Metropolis–Hastings algorithm for parameter inference in a nonlinear state space model.
  • Appendix H
    contains additional experiments in the Cauchy location model introduced in Section 2.4.
Code to reproduce the figures of this article can be found at: https://github.com/pierrejacob/unbiasedpoisson/.
Acknowledgements
Research of AL supported by EPSRC grants EP/R034710/1 and EP/Y028783/1. Research of DV supported by SERB (SPG/2021/001322).
Appendix A Unbiased MCMC with couplings
A.1 Couplings of MCMC algorithms
We recall some implementable coupling techniques for MCMC. We focus on Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithms (Hastings 1970). The goal is to couple generated trajectories such that exact meetings can occur. Relevant considerations can be found in Jacob et al. (2020), Wang et al. (2021), Papp & Sherlock (2024).
@@ A.1 Couplings of MCMC algorithms
When $\mu$ and $\nu$ are Normal distributions with the same variance, an alternative maximal coupling procedure is described in Algorithm A.3; it was proposed in Bou-Rabee et al. (2020). Its cost is deterministic and independent of $\mu$ and $\nu$. In the algorithmic description, $\varphi$ refers to the probability density function of the standard Normal distribution. We use Algorithm A.3 to couple the Metropolis-adjusted Langevin algorithm in Section 5.2 and Appendix E.
@@
A.2 Unbiased estimators of expectations and signed measures
$H_{k}=h(X_{k})+\sum_{j=1}^{\infty}\{h(X_{k+jL})-h(Y_{k+(j-1)L})\}.$ (A.1)
$H_{k:\ell}$ $=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}H_{t}$
$v_{t}=|\{s\in\{k,\ldots,\ell\}:\exists j\geq 1:s+jL=t\}|.$
We can re-write this as
$v_{t}=|\{n\in\{t-\ell,\ldots,t-k\}:\frac{n}{L}\in\mathbb{Z}^{+}\}|,$
$v_{t}=\lfloor(t-k)/L\rfloor-\lceil\max(L,t-\ell)/L\rceil+1.$ (A.2)
$H_{k:\ell}=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}h(X_{t})+\text{BC}_{k:\ell},$ (A.3)
where $\text{BC}_{k:\ell}$ refers to a “bias cancellation” term,
$\text{BC}_{k:\ell}=\sum_{t=k+L}^{\tau-1}\frac{v_{t}}{\ell-k+1}\left\{h(X_{t})-h(Y_{t-L})\right\}.$ (A.4)
$\hat{\pi}({\rm d}x)=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}\delta_{X_{t}}({\rm d}x)+\sum_{t=k+L}^{\tau-1}\frac{v_{t}}{\ell-k+1}\left\{\delta_{X_{t}}-\delta_{Y_{t-L}}\right\}({\rm d}x),$ (A.5)
as an unbiased approximation of $\pi$.
If we count the cost of sampling from the kernel $P$ as one unit, and the cost of sampling from $\bar{P}$ as one unit if the chains have met already, and two units otherwise, then the random cost of obtaining (A.3), or (A.5), equals $\max(L,\ell+L-\tau)+2(\tau-L)$ units.
A.3 Upper bounds on the distance to stationarity
$|\pi_{t}-\pi|_{\text{TV}}\leq\mathbb{E}\left[\max\left(0,\left\lceil\frac{\tau-L-t}{L}\right\rceil\right)\right].$ (A.6)
@@ B.1 Assumption on the meeting time
Applying Proposition B.1 below to $\psi(v)=\phi(v)=\vartheta v^{\alpha}$, there exists a constant $\beta$ such that for all $x,y\in{\mathcal{X}}$ and all $n\in{\mathbb{N}}$,
@@
In particular, with this $\phi$ and any $\sigma\in(0,1)$, denoting $H_{\phi}(t)=\int_{1}^{t}\frac{1}{\phi(u)}{\rm d}u$, there exists $c_{\sigma}>0$ such that $H_{\phi}^{-1}(\sigma n)\geq c_{\sigma}H_{\phi}^{-1}(n)$ and concavity and non-negativity of $\phi$ allows us to conclude that $\phi\circ H_{\phi}^{-1}(\sigma n)\geq c_{\sigma}\,\phi\circ H_{\phi}^{-1}(n)$. Finally, there exists $c^{\prime}_{\vartheta,\alpha}>0$ such that $\phi\circ H_{\phi}^{-1}(n)\geq c^{\prime}_{\vartheta,\alpha}(1+n)^{\frac{\alpha}{1-\alpha}}$. According to Proposition 4.3.2 in Douc et al. (2018), the drift condition $PV(x)\leq V(x)-\vartheta V^{\alpha}(x)+b\mathds{1}(x\in\mathcal{C})$ implies $\pi(V^{\alpha})<\infty$, which allows to apply Proposition 8 and the proof is concluded. ∎
Proposition B.1.
@@
$\int\pi({\rm d}x)\pi({\rm d}y){\mathbb{P}}_{x,y}(\tau>n)\leq\frac{\beta_{0}}{\phi\circ H_{\phi}^{-1}(\sigma n)},$
@@
Proof of Proposition B.1.
In this proof, for any function $\phi$, we use the notation $H_{\phi}(t)=\int_{1}^{t}\frac{1}{\phi(u)}{\rm d}u$ and $r_{\phi}(t)=\phi\circ H_{\phi}^{-1}(t)$. By Lemma 19.5.3 in Douc et al. (2018), for any $\sigma\in(0,1-b/d)$, $\bar{P}\bar{V}\leq\bar{V}-\bar{\phi}\circ\bar{V}+\bar{b}\mathds{1}_{\bar{\mathcal{C}}}$ where $\bar{\phi}=\sigma\phi$, $\bar{b}=2b$, $\bar{V}(x,y)=V(x)+V(y)-1$ and $\bar{\mathcal{C}}={\mathcal{C}}\times{\mathcal{C}}$.
@@
where $b^{\prime}=\bar{b}r_{\bar{\phi}}(1)/r_{\bar{\phi}}^{2}(0)$. Set $r^{0}_{\bar{\phi}}(n)=\sum_{j=0}^{n}r_{\bar{\phi}}(j)$ and define $M=\sup_{k}-r_{\bar{\phi}}^{0}(k-1)+b^{\prime}r_{\bar{\phi}}(k)/\epsilon$, which is finite since $r_{\bar{\phi}}(k)/r^{0}_{\bar{\phi}}(k-1)\to 0$ as $k$ tends to infinity.
@@
where $\bar{\Delta}(x,y)=\mathds{1}(x\neq y)$. The coupling kernel $\bar{P}$ is assumed to be sticky, i.e. $\bar{P}\bar{\Delta}(x,y)=0$ if $\bar{\Delta}(x,y)=0$. Hence, if $x=y$, the l.h.s. and r.h.s of (B.2) are null. We now assume that $x\neq y$. Using $\bar{\Delta}\leq 1$, we get
@@
where we have used Assumption 10 and the fact that $b^{\prime}r_{\bar{\phi}}(k)-\epsilon(r^{0}_{\bar{\phi}}(k-1)+M)\leq 0$ by definition of $M$. Since $x\neq y$, we have $\bar{\Delta}(x,y)=1$ and hence $\bar{P}[\bar{\Delta}S_{k+1}](x,y)\leq[\bar{\Delta}S_{k}](x,y)$. At this point, (B.2) is proved for all $x,y\in{\mathcal{C}}$. Now write
@@
$A_{n}=\left\{(x,y)\in\mathbb{X}\times\mathbb{X}:\;V(x)+V(y)>H^{-1}_{\bar{\phi}}(n)\right\},$
@@
Proposition B.2 (Douc et al. 2018, Proposition 21.2.3).
@@
and $P$ is aperiodic, and ergodic of degree 2.
@@
and since Assumption 2 provides $\kappa>1$ and $\mathbb{E}_{x,\pi}[\tau^{\kappa}]<\infty$ for $\pi$-almost all $x$, $\left\|P^{t}(x,\cdot)-\pi\right\|_{{\rm TV}}$ goes to zero as $t\to\infty$ for $\pi$-almost all $x$ and so $P$ is aperiodic by Douc et al. (2018, Lemma 9.3.9). We also observe that (B.4) implies (B.3) and since $\kappa>1$ we have
@@
$\leq\mathbb{E}_{x,\pi}\left[\mathds{1}(\tau>t)\left\{\left|h(X_{t})\right|+\left|h(Y_{t})\right|\right\}\right],$
@@
$\left|P^{t}h(x)\right|^{p}$ $\leq\mathbb{E}_{x,\pi}\left[\mathds{1}(\tau>t)\left\{\left|h(X_{t})\right|+\left|h(Y_{t})\right|\right\}\right]^{p}$
$\leq\mathbb{E}_{x,\pi}\left[\mathds{1}(\tau>t)\left\{\left|h(X_{t})\right|+\left|h(Y_{t})\right|\right\}^{p}\right]$
$\leq 2^{p-1}\mathbb{E}_{x,\pi}\left[\mathds{1}(\tau>t)\left\{\left|h(X_{t})\right|^{p}+\left|h(Y_{t})\right|^{p}\right\}\right].$
@@
$\leq 2^{p-1}\mathbb{E}_{\pi\otimes\pi}\left[\mathds{1}(\tau>t)\left\{\left|h(X_{t})\right|^{p}+\left|h(Y_{t})\right|^{p}\right\}\right]$
@@
and we conclude by appealing to Proposition B.2. ∎
Remark B.5.